# GSCache: Real-Time Radiance Caching for Volume Path Tracing using 3D Gaussian Splatting

### v1

@@ 1 Introduction
In this work, we introduce a real-time path-space cache for volume rendering that stores attenuated radiance at different path lengths and can quickly adapt to changes in the scene, such as lighting, transfer function, or slicing operations. The cache is represented as a multi-level collection of 3D Gaussians. Each level caches the attenuated path-space radiance for paths of a specific length. Cache sizes are sub-sampled from higher to lower levels, similar to the construction of MIP levels in texture processing. This is done since paths of length $n$ generally contribute more to the final image than paths of length $m$, where $m>n$. This property allows us to allocate more compute resources and representational power to paths that have a higher image contribution. The cache is optimized in real time with noisy samples from the renderer. The only information needed from the rendering application is the attenuated radiance value and the path length that produced it.
@@
With these contributions, we hope to further the design and quality of scientific renderers and encourage the use of radiance caching in scientific applications and beyond.
@@ 2.1 Radiance Caching
Radiance caching aims to trade variance for bias to improve the image quality of path tracing applications. In the context of path tracing, variance is the result of sampling decisions along a path while bias is generally the result of making assumptions about radiance distributions in a scene that are not analytically tractable. In most cases (ir-)radiance is cached by storing radiance probes in world-accessible spatial data structures that can be sampled during rendering [19, 10, 39]. Recently, some approaches have made use of deep learning to store and access radiance information [29, 2, 8]. Early approaches usually focus on caching indirect illumination after the first bounce. This means that all paths that terminate into the cache do so after a single bounce. A downside of such approaches is that using the cache so soon can introduce considerable error from bias. This is why later approaches try to mask the bias by allowing longer paths [14, 29]. There has also been work specifically targeting volume visualization. Khlebnikov et al. [16] introduce a parallel GPU-bases irradiance caching approach for interactive volume rendering. Šmajdek et al. [41] introduce radiance caching to address mixed-modality scenarios where isosurface and volume rendering are combined.
One crucial consideration to make when allowing for different path lengths is to decide when to terminate a path and instead to read stored radiance from the cache and update its entries. Throughout the literature, several heuristics have been developed to address this issue. For example, Müller et al. [29] terminate paths based on their spatial spread using a heuristic proposed by Baekert et al. [3]. Kandlbinder et al. [13] develop a variance-based heuristic that offers a better control of the bias error bound when sampling the cache.
@@ 2.2 Scientific Volume Rendering
Volume rendering is vital to the field of scientific visualization. Broadly speaking, applications of scientific rendering can be categorized into two groups—visualization for analysis and visualization for illustration. The former group relies on rendering techniques that offer interactive framerates and user interactions, such as cutting planes and transfer function editors. We commonly see direct volume rendering (DVR) as the rendering method of choice in this group. Illustrative visualization, on the other hand, prioritizes image quality over interactivity and customization. It is closer, in purpose, to more traditional offline rendering scenarios. In this case, path tracing with physically-based illumination models is preferable. Both types of rendering and the underlying optical models for volume visualization have been aptly summarized by Max et al. [28]. Despite promises of realism, calculating physically-based volumetric illumination via the radiative transfer equation (RTE) is expensive. For a long time, research has focused on finding appropriate approximations to enable real-time rendering of scientific datasets. Jönsson et al. [12] provide an excellent summary of this category of techniques. In the following, we highlight some notable works in this area. Gradient-based methods were one of the first ways to render volume datasets. Levoy et al. [23] use volume gradients to apply surface shading models like the Blinn-Phong reflection model. Local shading is a simple step up from gradient-based rendering. Ambient occlusion techniques [36, 37, 38] have been popular in the domain as they are relatively easy to implement and grant viewers a greater sense of depth, which makes it easier to discern details in complex or abstract volume datasets. Next, slice-based methods like Kniss et al.’s half-angle slicing [17, 18] provide a more sophisticated model and were one of the first to simulate multi-scatter illumination. Lastly, physically-based particle tracking [47] and path tracing allow for high levels of realism with full global illumination. Here, Zhang et al. [50] use pre-computed photon maps while others [5, 20] work on ray tracing based approaches. Jönsson et al. [11] expand on the idea of photon mapping for scientific visualization and introduce an approach that selectively recomputes photon traces based on transfer function changes. Lastly, Bauer et al. [2] transform photon maps into implicit neural representations, which improves rendering speeds, enabling interactive visualization. Our work is related to these techniques in that we aim to provide higher performance and lower variance for scientific path tracing applications.
@@ 2.3 Machine Learning for Scientific Visualization
Scientific visualization research has been comparatively slow to adapt machine learning and deep learning into their repertoire. Nevertheless, recent years have witnessed several interesting applications of machine learning techniques to address pervasive problems in scientific visualization and rendering.
@@
Lastly, storage poses additional challenges over traditional rendering applications. Volume datasets tend to be large in size with multiple parameter grids and even multiple volumes composed as a time series. Neural networks have proved to be a powerful tool to implicitly represent and compress such volumes [9, 43, 49]. Weiss et al. [43] and Wu et al. [49] develop implicit neural representations for volume compression to address this issue. More recently, Wu et al. [48] have extended their neural representation using hypernetworks for temporal volume compression.
@@ 3.2 Cache Initialization
$s_{i}=\frac{((\mu_{N}+2*\sigma_{N})\land\frac{1}{3}\sum_{j=0}^{2}{d_{ij}})}{2}$ (2)
@@ 3.3 Path Termination Heuristics
The overall goal of such a path termination heuristic is to determine when the benefits of tracing a full, unbiased path outweigh the benefits of using a biased cache sample instead. Traditionally, an effective measure to determine this has been to estimate the current path’s contribution if it were to leave the volume after the current path vertex to receive radiance from the environment outside the volume. In the case of basic uniform path tracing, this estimation occurs when the path naturally terminates, which happens when the tracking algorithm does not find a valid interaction with the volume indicating that the path has left the volume bounds. This approach is easy to implement but suffers from the fact that paths have to be traced to their natural conclusion, foregoing the potential savings of terminating a low contribution ray early. To improve upon the naïve approach, we leverage the commonly implemented technique of next-event-estimation (NEE). This technique estimates direct light contribution at every path vertex. The light sample is then combined with the continued path contribution using multiple importance sampling. Having direct light samples at every path vertex allows us to continuously estimate a path’s contribution as we trace it through the volume. In our implementation, we distinguish two cases—natural and early path termination—and we propose two simple cache sampling strategies for both cases.
@@ 3.3.1 Natural Path Termination
The base case in most volume path tracers is reached when there is no new valid interaction with the volume, and the path naturally leaves the volume. Since path termination is not deliberately sampled but rather an unforeseeable outcome of phase function sampling, terminating directions are chosen according to the volume’s scattering behavior. The majority of such terminal path segments typically does not intersect with any of the light sources in the scene, leading to zero radiance contribution along the ray, effectively wasting the whole path sample. In scenes with infinite area light sources, samples might contribute a small amount of radiance. We can make use of this fact to sample our cache. If the path terminates with non-zero radiance contribution, we use the unbiased sample; otherwise, we determine the cached radiance at the terminating path length and use this value instead. In this way, every path that terminates contributes a meaningful amount of radiance to the final image.
@@ 3.3.2 Early Path Termination
In addition to sampling the cache for naturally terminated paths, we also consider terminating paths early into the cache. This has the benefit that we not only get a less noisy sample but also do so at a performance benefit as reading the cache is cheaper than sampling the path to its natural termination.
The heuristic uses the luminance of the current path throughput $Tr$ to generate a termination probability $p$ (see Algorithm 1). The Algorithm takes a list of path vertex albedos $\sigma_{i}$ to compute $Tr$. At the same time we keep track of $\beta_{i}$ to importance sample cache hits. The details of the algorithm are covered in the subsequent paragraphs. The heuristic is similar to the Russian Roulette termination technique that is commonly used in path tracers to cut long paths with little contribution short. Unlike Russian Roulette, our method uses a much higher initial throughput threshold to become active. In our tests, we consider any throughput values less than 0.90.90.9. The reason for having this threshold is to guarantee that a fraction of paths will continue regardless of any cache sampling probability to ensure that we can gather paths of all lengths. Additionally, we employ a user-defined termination coefficient $C$, which modifies the baseline probability in favor or against the current throughput. The final probability is used to determine if the current path should be terminated into the cache and, if not, is used to importance sample the continuation of the path (see Section 3.4).
@@ 3.4 Cache Sampling
If a path is terminated at depth $n$, we want to determine if the final radiance should come from the cache or the path’s contribution until vertex $n$. To make this decision, we look to the current total path throughput $Tr$ as an indicator of anticipated path contribution (Figure 2). We sample the radiance cache based on the anticipated path contribution to select either the value in the cache at the corresponding level or use the current unbiased radiance that we accumulated through next-event-estimation.
@@ 3.4.2 Cascaded Cache Radiance Importance Sampling
Adjusting path throughput $Tr_{out}$ by the probability of a cache hit is not enough to ensure correct cache sampling. Another issue arises when there is a cache hit. The cached values are pre-attenuated and directly represent the radiance at path length $n$. However, this value does not take into account the possibilities of early terminations at length $i\in[1,n-1]$ (Figure 3). Since paths could have terminated at any path vertex before the current one, only a fraction of paths reach length $n$. We need to account for this fact by keeping track of the product of all prior cache sampling probabilities $Tr$ and adjusting the cached radiance by this product (see Algorithm 1). Not doing so will result in images that are generally darker than the unbiased reference. To keep track of the product path sampling probability $\beta$, we initialize it to 111 and multiply it with $Tr$ any time the cache is missed. When the cache is hit at depth $n$, we divide the cached radiance by $\beta_{n-1}$ to account for the fact that the cache was missed $n-1$ times. The effective attenuation applied to a path sample is as follows.
@@ 3.5 Cache Training
At the end of one sample pass, each pixel will have at most one valid entry across the intermediate path buffers. At this time, the samples are attenuated, meaning that their total path throughput is taken care of. However, these samples are ultimately noisy as they represent the radiance from only one possible path of length $n$ through the volume. Conventionally, Gaussian splatting applications use clean target images such as photographs to calculate gradients for optimization. This poses a potential problem for our application.
We make the important observation that the expected value of samples in the buffers behaves analogously to regular path tracing where paths of all lengths are collected indiscriminately. This means that if we were to take an infinite number of samples, we could accumulate a clean image for each of the buffers representing the radiance contribution in path subspace $n$.
@@ 3.6 Optimization and Hyperparameters
Densification and Pruning. We do not use the original densification and pruning strategies introduced by Kerbl et al. [15]. The noise of our training samples makes the splitting, cloning, and pruning mechanisms too unstable for real time training. Furthermore, removing these heuristic-based strategies improved not only stability but also runtime of our method.
Loss. The data we are fitting has a high dynamic range (HDR) with potentially unbounded radiance values. This is in contrast to most scene representation tasks, which will draw samples from LDR sources like photographs. Commonly used loss terms like the L1 loss are ill-fit to adapt to HDR data as very bright samples generate disproportionately large gradients, which can overshadow the contribution of other samples. To address this, we adapt the HDR loss proposed by Lehtinen et al. [22], which contains a normalization term and reads as follows.
@@ 3.7 Implementation
We implement the radiance cache using C++ and CUDA. Optimization is handled via the PyTorch C++ API [7] and is directly integrated with the C++ runtime. We expose the caching library via a C-style API, which allows it to be easily adapted to various applications and languages. Furthermore, we provide Python bindings for the API via pybind11 [46].
@@ 4.2 Datasets
Our method is evaluated on a variety of volume datasets. In the following, we describe each dataset’s characteristics (see Table 1). We chose the datasets for this study to show a representative selection of data from medicine, biology, physics, and other fields. Since the present study concerns real-time applications, we did not consider extreme-scale datasets, which would require specialized data streaming and rendering techniques and might limit the system to offline rendering, weakening the justification for a cache. For each of the datasets we use a predefined 1D transfer function which remains unchanged throughout the evaluation.
@@ 4.3 Visual Quality
In Figure 5, we compare the image quality of renderings from several different scenes. The results show that our method produces significantly higher image quality at low sample rates. Images are less noisy, as evidenced by the generally higher PSNR, and overall image quality notably exceeds that of baseline renderings and state-of-the art caching [29] results at equivalent sampling rates. We highlight the cache quality over time in Figure 6. After a cold start, it takes less than 16 samples for the cache to adapt to a point where it outperforms the baseline renderer in terms of visual quality. As rendering progresses, we significantly improve over the baseline and achieve superior image quality. Lastly, we investigate the effects of different cache sampling coefficients $C$, which we define in Section 3.3.2. The results show how $C$ can be used to move along the bias-variance scale to find a suitable balance between image quality, rendering speed, and unbiasedness (Figure 7). In practice, the lower bound $C=0.0$ is likely not going to be used as it corresponds to pure cache sampling without introducing unbiased paths, which implicitly precludes the possibility of cache training. We found a value of $C=0.5$ to be an adequate setting in our test scenes. Furthermore, we discuss the implications of different $C$ values for rendering performance in the next section.
Table 2: Frame timings of our method on different datasets captured from screen-filling camera fly-through of 200 frames. We allowed for a 40-frame warm-up period to initialize the cache and show path tracing time (PT) for all four methods and splatting time (ST) , inference time (IT), and optimization time (OT) for our method and NRC [29]. IT includes a composition pass to back-propagate and compose cached radiance. The notation $PT_{nee*}$ denotes NEE with the addition of our path termination heuristic and $PT_{\text{nee}**}$ additionally adds path records and training sample collection for NRC [29]. All values were captured at $C=0.5$ with $N=300$k initial points. All timings are in milliseconds.
@@
Figure 6: Comparison of cache state after a cold-start training for 64 frames. Frame 1 shows the freshly-initialized state with raw albedo colors. Over the next few frames, the cache quickly adapts to the scene. We show the cache states after 1, 4, 16, and 64 frames and juxtapose them with an uncached frame capture. Images are rendered at 16 SPP, and quality metrics are computed compared to the unbiased reference image. Note that at 16 SPP, the cache provides diminishing returns as more unbiased paths are accumulated. Nevertheless, our cache outperforms the unbiased renderer at the same number of samples.
@@
Figure 9: Cache amortization behavior example from the FullBody dataset. NEE* denotes NEE times with our ray termination heuristic. At low sample counts, our method performs better than the baseline NEE path tracer. As SPP increase, the time to generate the cache buffers stays constant, and the overall runtime for our method grows slower than both baseline approaches.

### v2

@@ 1 Introduction
In this work, we introduce a real-time path-space cache for volume rendering that stores attenuated radiance at different path lengths and can quickly adapt to changes in the scene, such as lighting, transfer function, or slicing operations. The cache is represented as a multi-level collection of 3D Gaussians. Each level caches the attenuated path-space radiance for paths of a specific length. Cache sizes are sub-sampled from higher to lower levels, similar to the construction of MIP levels in texture processing. This is done since paths of length $n$ generally contribute more to the final image than paths of length $m$, where $m>n$. This property allows us to allocate more computational resources and representational power to paths that have a higher image contribution. The cache is optimized in real time with noisy samples from the renderer. The only information needed from the rendering application is the attenuated radiance value and the path length that produced it.
@@
With these contributions, we hope to support the design and quality of scientific renderers and encourage the use of radiance caching in scientific applications and beyond.
@@ 2.1 Radiance Caching
Radiance caching aims to trade variance for bias to improve the image quality of path tracing applications. In the context of path tracing, variance is the result of sampling decisions along a path, while bias is generally the result of making assumptions about radiance distributions in a scene that are not analytically tractable. In most cases (ir-)radiance is cached by storing radiance probes in world-accessible spatial data structures that can be sampled during rendering [23, 13, 44]. Recently, some approaches have made use of deep learning to store and access radiance information [34, 2, 10]. Early approaches usually focus on caching indirect illumination after the first bounce. This means that all paths that terminate into the cache do so after a single bounce. A downside of such approaches is that using the cache so soon can introduce considerable error from bias. This is why other approaches try to mask the bias by allowing longer paths [17, 34]. There has also been work specifically targeting volume rendering. Khlebnikov et al. [19] introduce a parallel GPU-based irradiance caching approach for interactive volume rendering. Jakob et al. [12] use hierarchical Gaussian mixture models to derive scene radiance from photon traces and Šmajdek et al. [46] introduce radiance caching to address mixed-modality scenarios where isosurface and volume rendering are combined.
One crucial consideration to make when allowing for different path lengths is to decide when to terminate a path and instead to read stored radiance from the cache and update its entries. Throughout the literature, several heuristics have been developed to address this issue. For example, Müller et al. [34] terminate paths based on their spatial spread using a heuristic proposed by Baekert et al. [3]. Kandlbinder et al. [16] develop a variance-based heuristic that offers better control of the bias error bound when sampling the cache.
@@ 2.2 Scientific Volume Rendering
Volume rendering is vital to the field of scientific visualization. Broadly speaking, applications of scientific rendering can be categorized into two groups—visualization for analysis and visualization for illustration. The former group relies on rendering techniques that offer interactive framerates and user interactions, such as cutting planes and transfer function editors. We commonly see direct volume rendering (DVR) as the rendering method of choice in this group. Illustrative visualization, on the other hand, prioritizes image quality over interactivity and customization. It is closer, in purpose, to more traditional offline rendering scenarios. In this case, path tracing with physically-based illumination models is preferable. Both types of rendering and the underlying optical models for volume visualization have been aptly summarized by Max et al. [32]. Despite promises of realism, calculating physically-based volumetric illumination via the radiative transfer equation (RTE) is expensive. For a long time, research has focused on finding appropriate approximations to enable real-time rendering of scientific datasets. Jönsson et al. [15] provide an excellent summary of this category of techniques. In the following, we highlight some notable works in this area. Gradient-based methods were one of the first ways to render volume datasets. Levoy et al. [27] use volume gradients to apply surface shading models like the Blinn-Phong reflection model. Local shading is a simple step up from gradient-based rendering. Ambient occlusion techniques [41, 42, 43] have been popular in the domain as they are relatively easy to implement and grant viewers a greater sense of depth, which makes it easier to discern details in complex or abstract volume datasets. Next, slice-based methods like Kniss et al.’s half-angle slicing [21, 22] provide a more sophisticated model and were one of the first to simulate multi-scatter illumination. Lastly, physically-based particle tracking [51] and path tracing allow for high levels of realism with full global illumination. Here, Zhang et al. [56] use pre-computed photon maps while others [6, 24] work on ray tracing-based approaches. Jönsson et al. [14] expand on the idea of photon mapping for scientific visualization and introduce an approach that selectively recomputes photon traces based on transfer function changes. Lastly, Bauer et al. [2] transform photon maps into implicit neural representations, which improves rendering speeds, enabling interactive visualization. Our work is related to these techniques in that we aim to provide higher performance and lower variance for scientific path tracing applications.
@@ 2.3 Machine Learning for Scientific Visualization
Scientific visualization research has been comparatively slow to adapt machine learning and deep learning into its repertoire. Nevertheless, recent years have witnessed several interesting applications of machine learning techniques to address pervasive problems in scientific visualization and rendering.
@@
Lastly, storage poses additional challenges over traditional rendering applications. Volume datasets tend to be large, with multiple parameter grids and even multiple volumes composed as a time series. Neural networks have proved to be a powerful tool to implicitly represent and compress such volumes [11, 48, 53]. Weiss et al. [48] and Wu et al. [53] develop implicit neural representations (INR) for volume compression to address this issue. Zavorotny et al. [55] address INR rendering bottlenecks while, more recently, Wu et al. [52] and Gadirov [9] have extended neural visualization methods using hypernetworks for parameter-space exploration.
@@ 3.2 Cache Initialization
$s_{i}=\frac{(\mu_{N}+2*\sigma_{N})\land\frac{1}{3}\sum_{j=0}^{2}{d_{ij}}}{2}$ (2)
@@ 3.3 Path Termination Heuristics
The overall goal of such a path termination heuristic is to determine when the benefits of tracing a full, unbiased path outweigh the benefits of using a biased cache sample instead. Traditionally, an effective measure to determine this has been to estimate the current path’s contribution if it were to leave the volume after the current path vertex to receive radiance from the environment outside the volume. In the case of basic uniform path tracing, this estimation occurs when the path naturally terminates, which happens when the tracking algorithm does not find a valid interaction with the volume, indicating that the path has left the volume bounds. This approach is easy to implement but suffers from the fact that paths have to be traced to their natural conclusion, foregoing the potential savings of terminating a low-contribution ray early. To improve upon the naïve approach, we leverage the commonly implemented technique of next-event-estimation (NEE). This technique estimates direct light contribution at every path vertex. The light sample is then combined with the continued path contribution using multiple importance sampling. Having direct light samples at every path vertex allows us to continuously estimate a path’s contribution as we trace it through the volume. In our implementation, we distinguish two cases—natural and early path termination—and we propose two simple cache sampling strategies for both cases.
@@ 3.3.1 Natural Path Termination
The base case in most volume path tracers is reached when there is no new valid interaction with the volume, and the path naturally leaves the volume. Since path termination is not deliberately sampled but rather an unforeseeable outcome of phase function sampling, terminating directions are chosen according to the volume’s scattering behavior. The majority of such terminal path segments typically does not intersect with any of the light sources in the scene, leading to zero radiance contribution along the ray, effectively wasting the whole path sample. In scenes with infinite area light sources, samples might contribute a small amount of radiance. We can make use of this fact to sample our cache. If the path terminates with a non-zero radiance contribution, we use the unbiased sample; otherwise, we determine the cached radiance at the terminating path length and use this value instead. In this way, every path that terminates contributes a meaningful amount of radiance to the final image.
@@ 3.3.2 Early Path Termination
In addition to sampling the cache for naturally terminated paths, we also consider terminating paths early into the cache. This has the benefit that we not only get a less noisy sample but also do so at a performance gain, as reading the cache is cheaper than sampling the path to its natural termination.
The heuristic uses the luminance of the current path throughput $Tr$ to generate a termination probability $p$ (see Algorithm 1). The Algorithm takes a list of path vertex albedos $\sigma_{i}$ to compute $Tr$. At the same time, we keep track of $\beta_{i}$ to importance sample cache hits. The details of the algorithm are covered in the subsequent paragraphs. The heuristic is similar to the Russian Roulette termination technique that is commonly used in path tracers to cut long paths with little contribution short. Unlike Russian Roulette, our method uses a much higher initial throughput threshold to become active. In our tests, we consider any throughput values less than 0.90.90.9. The reason for having this threshold is to guarantee that a fraction of paths will continue regardless of any cache sampling probability to ensure that we can gather paths of all lengths. Additionally, we employ a user-defined termination coefficient $C$, which modifies the baseline probability in favor or against the current throughput. The final probability is used to determine if the current path should be terminated into the cache and, if not, is used to importance sample the continuation of the path (see Section 3.4).
@@ 3.4 Cache Sampling
If a path is terminated at depth $n$, we want to determine if the final radiance should come from the cache or the path’s contribution until vertex $n$. To make this decision, we look to the current total path throughput $Tr$ as an indicator of anticipated path contribution (Figure 2). We sample the radiance cache based on the anticipated path contribution to select either the value in the cache at the corresponding level or use the current unbiased radiance that we accumulated through next-event estimation.
@@ 3.4.2 Cascaded Cache Radiance Importance Sampling
Adjusting path throughput $Tr_{out}$ by the probability of a cache hit is not enough to ensure correct cache sampling. Another issue arises when there is a cache hit. The cached values are pre-attenuated and directly represent the radiance at path length $n$. However, this value does not take into account the possibilities of early terminations at length $i\in[1,n-1]$ (Figure 3). Since paths could have terminated at any path vertex before the current one, only a fraction of paths reach length $n$. We need to account for this fact by keeping track of the product of all prior cache sampling probabilities $Tr$ and adjusting the cached radiance by this product (see Algorithm 1). Not doing so will result in images that are generally darker than the unbiased reference. To keep track of the product path sampling probability $\beta$, we initialize it to 111 and multiply it by $Tr$ any time the cache is missed. When the cache is hit at depth $n$, we divide the cached radiance by $\beta_{n-1}$ to account for the fact that the cache was missed $n-1$ times. The effective attenuation applied to a path sample is as follows.
@@ 3.5 Cache Training
At the end of one sample pass, each pixel will have at most one valid entry across the intermediate path buffers. At this time, the samples are attenuated, meaning that their total path throughput is taken care of. However, these samples are ultimately noisy as they represent the radiance from only one possible path of length $n$ through the volume. Conventionally, Gaussian splatting applications use clean target images, such as photographs, to calculate gradients for optimization. This poses a potential problem for our application.
We make the important observation that the expected value of samples in the buffers behaves analogously to regular path tracing, where paths of all lengths are collected indiscriminately. This means that if we were to take an infinite number of samples, we could accumulate a clean image for each of the buffers representing the radiance contribution in path subspace $n$.
@@ 3.6 Optimization and Hyperparameters
Densification and Pruning. We do not use the original densification and pruning strategies introduced by Kerbl et al. [18]. The noise of our training samples makes the splitting, cloning, and pruning mechanisms too unstable for real-time training. Furthermore, removing these heuristic-based strategies improved not only stability but also the runtime of our method.
Loss. The data we are fitting has a high dynamic range (HDR) with potentially unbounded radiance values. This is in contrast to most scene representation tasks, which will draw samples from LDR sources like photographs. Commonly used loss terms like the L1 loss are ill-suited to adapt to HDR data as very bright samples generate disproportionately large gradients, which can overshadow the contribution of other samples. To address this, we adapt the HDR loss proposed by Lehtinen et al. [26], which contains a normalization term and reads as follows.
@@ 3.7 Implementation
We implement the radiance cache using C++ and CUDA. Optimization is handled via the PyTorch C++ API [8] and is directly integrated with the C++ runtime. We expose the caching library via a C-style API, which allows it to be easily adapted to various applications and languages.
@@ 4.2 Datasets
Our method is evaluated on a variety of volume datasets. In the following, we describe each dataset’s characteristics (see Table 1). We chose the datasets for this study to show a representative selection of data from medicine, biology, physics, and other fields. Since the present study concerns real-time applications, we did not consider extreme-scale datasets, which would require specialized data streaming and rendering techniques and might limit the system to offline rendering, weakening the justification for a cache. For each of the datasets, we use a predefined 1D transfer function which remains unchanged throughout the evaluation.
@@ 4.3 Visual Quality
In Figure 5, we compare the image quality of renderings from several different scenes. The results show that our method produces significantly higher image quality at low sample rates. Images are less noisy, as evidenced by the generally higher PSNR, and overall image quality notably exceeds that of baseline renderings and state-of-the-art caching [34] results at equivalent sampling rates. We highlight the cache quality over time in Figure 6. After a cold start, it takes less than 16 samples for the cache to adapt to a point where it outperforms the baseline renderer in terms of visual quality. As rendering progresses, we significantly improve over the baseline and achieve superior image quality. Lastly, we investigate the effects of different cache sampling coefficients $C$, which we define in Section 3.3.2. The results show how $C$ can be used to move along the bias-variance scale to find a suitable balance between image quality, rendering speed, and unbiasedness (Figure 7). In practice, the lower bound $C=0.0$ is likely not going to be used as it corresponds to pure cache sampling without introducing unbiased paths, which implicitly precludes the possibility of cache training. We found a value of $C=0.5$ to be an adequate setting in our test scenes. Furthermore, we discuss the implications of different $C$ values for rendering performance in the next section.
Table 2: Frame timings of our method on different datasets captured from a screen-filling camera fly-through of 200 frames. We allowed for a 40-frame warm-up period to initialize the cache and show path tracing time (PT) for all four methods and splatting time (ST) , inference time (IT), and optimization time (OT) for our method and NRC [34]. IT includes a composition pass to back-propagate and compose cached radiance. The notation $PT_{nee*}$ denotes NEE with the addition of our path termination heuristic and $PT_{\text{nee}**}$ additionally adds path records and training sample collection for NRC [34]. All values were captured at $C=0.5$ with $N=300$k initial points. All timings are in milliseconds.
@@
Figure 6: Comparison of cache state after a cold-start training for 64 frames. Frame 1 shows the freshly initialized state with raw albedo colors. Over the next few frames, the cache quickly adapts to the scene. We show the cache states after 1, 4, 16, and 64 frames and juxtapose them with an uncached frame capture. Images are rendered at 16 SPP, and quality metrics are computed compared to the unbiased reference image. Note that at 16 SPP, the cache provides diminishing returns as more unbiased paths are accumulated. Nevertheless, our cache outperforms the unbiased renderer at the same number of samples.
@@
Figure 9: Cache amortization behavior example from the FullBody dataset. NEE* denotes NEE times with our ray termination heuristic. At low sample counts, our method performs better than the baseline NEE path tracer. As SPP increase, the time to generate the cache buffers stays constant, and the overall runtime for our method grows more slowly than both baseline approaches.